If I have a 55.80% as a grade right now, how many more tests will be needed to be taken and what scores would they need to be to get a 70%?
Answer by LisaJ(11)

It depends on how many tests you have already taken. You need to take the average of all your test scores.
For example, if you have taken one test so far, you can take one more test and get 100%.
(100 + 55.8)2 = 77.9 %
If you get 90% on the second test:
(90 + 55.8)/2 = 72.9 %
If you have taken 4 tests so far, you can take 4 more tests, and get 90% on each.
(4*55.8 + 4*90)/8 = 72.9%
Or take 3 more tests, and get 100% on them:
(4*55.8) + 3*100) /7 = 75%
Or take 5 more tests, and get 82% on them:
(4*55.8) + 5*82) /9 = 70%
Find out how many more test there will be, and then you can figure out how high you have to score on them.